OpenStudy (anonymous):

find the second derivative of the function. f(x)=(2x^4+1)^7

OpenStudy (campbell_st):

can you find the 1st derivative..?

OpenStudy (anonymous):

yea

OpenStudy (goformit100):

do you know chain rule of derivative

OpenStudy (campbell_st):

what is it...?

OpenStudy (anonymous):

let me work it out.....

OpenStudy (anonymous):

====7(2x^4+1)^6*8x

OpenStudy (anonymous):

u forgot 8x^exponent

OpenStudy (anonymous):

oooo 8x^3

OpenStudy (anonymous):

if we expand by binomail theorem it can be more easy to find second derivative

OpenStudy (campbell_st):

you are missing a power but its basically correct.. \[f'(x) = 56x^3(2x^4 + 1)^6\] do you know how to use the product rule... \[f(x) = g(x)\times h(x)\] where \[g(x) = 56x^3...... and ....h(x) = (2x^4 + 1)^6\]

OpenStudy (anonymous):

yes!

OpenStudy (campbell_st):

so does that help...?

OpenStudy (anonymous):

yes. the only problem is that my teacher gave me the answer to be =====56x^2(1+2x^4)^5(3+54x^4)

OpenStudy (anonymous):

will i get this if i apply the product rule?

OpenStudy (campbell_st):

ok... so you teacher has taken out common factors of 56x^2 and (2x^4 + 1)^5

OpenStudy (campbell_st):

can you try the product rule on the 1st derivative and then we'll look at the factoring

OpenStudy (anonymous):

im not sure how to do that because just looking at the problem i would think using chain rule

OpenStudy (campbell_st):

ok... \[g(x) = 56x^3 .....so ....g'(x) = 3 \times 56x^2\] and \[h(x) = (2x^4 + 1)^6 .... so.... h'(x) = 6 \times(2x^4 + 1)^5 \times 8x^3\] does this make sense..?

OpenStudy (anonymous):

this the chain rule agin right on the 2nd derivative?

OpenStudy (campbell_st):

then \[f''(x) = g(x)\times h'(x) + h(x) \times g'(x)\]

OpenStudy (anonymous):

ooooooooooooo chain rule on the 2nd derivative and then product rule?

OpenStudy (campbell_st):

not sure how your teacher got 54x^4 but we'll work through it

OpenStudy (anonymous):

ok :] no it was 56x^2

OpenStudy (campbell_st):

ok... lets see \[f"(x) = 56x^3 \times 48x^3(2x^4 + 1)^5 + (2x^4 + 1)^6 \times 3 \times 56x^2\] does that make sense...?

OpenStudy (campbell_st):

aand it all works out to your teacher's answer...

OpenStudy (anonymous):

i'll try this way on my own thanks so much! but okay doing it the other way, chainrule, then chain rule then product rule the answer is still correct right?

OpenStudy (campbell_st):

yep... so I'm going to rewrite the times a little \[56x^2 \times x \times 48x^3 \times(2x^4 + 1)^5 + (2x^4 + 1)^5 \times (2x^4 +1) \times 3 \times 56x^2\] which shows that the common factor is \[56x^2(2x^4 + 1)^5\]

OpenStudy (campbell_st):

this means that you will have \[56x^2(2x^4 + 1)^5 (x \times 48x^3 + (2x^4 + 1) \times 3)\] or \[56x^2(2x^4 + 1)^5(48x^4 + 6x^4 + 1)\] which finally simplifies to \[56x^2(2x^4 + 1)^5( 54x^4 + 3)\] you teacher's answer.

OpenStudy (anonymous):

TTTTTTTTTTHHHHHHHHHHHHAAAAAAAAAAAAAAAANNNNNNNNNNNNNNNNNKKKKKKKKKKKKKKKKKKKKKK YYYYYYYYYYYYYYYOOOOOOOOOOOOOOUUUUUUUUUUUUU!!!!!

OpenStudy (campbell_st):

and you need to use chain rule and product rule for the 2nd derivative.

OpenStudy (campbell_st):

opps slight typo... 2nd last line of the solution the brackets should read \[(48x^4 + 6x^4 + 3)\]